OpenStudy (anonymous):

The magnitude and direction of vectors u and v are given. Find vector w's polar coordinates.

OpenStudy (anonymous):

You could convert to rectangular coordinates, combine and convert back. For example in 1) Converting u we get x=136cos220 and y=136sin220 so rectangular coordinates are u=(-104.184,-87.419) and v=(197cos300,197sin300)=(98.5,-170.607) 2u-v=(-306.868,-4.231) magnitude=sqrt(306.868^2+4.231^2)=306.... tan(bearing)=y/x=-4.231/306.868) The reference angle is 0.79degrees, and since it is in the 3rd quadrant the bearing is 180.79 degrees.